OpenStudy (anonymous):

STATS help!!! A standard chemistry examination administered nationally by the American Chemical Society has a mean of 500 and a standard deviation of 90. What is the probability that the average of a random sample of examination scores of 25 students will be between 450 and 500?

OpenStudy (amistre64):

any thoughts?

OpenStudy (anonymous):

not sure :\

OpenStudy (amistre64):

well, persoanlly i would work a z score ... maybe a t score

OpenStudy (amistre64):

theres usually a flow chart that tells when to use which depending on conditions i believe this is a zscore tho

OpenStudy (amistre64):

do you ahve a ti83 to play with on this?

OpenStudy (anonymous):

@amistre64 nope

OpenStudy (amistre64):

then tables are your thing

OpenStudy (amistre64):

how many standard deviations is 450 from men of 500?

OpenStudy (anonymous):

0.56?

OpenStudy (amistre64):

50/90 = 5/9 = .56 is fair now a table z values should be in your material that we can use this with

OpenStudy (anonymous):

if I'm looking across the 0.6 z score, which column do i look down?

OpenStudy (amistre64):

row(0.5)+ col(.06)

OpenStudy (anonymous):

0.7123

OpenStudy (amistre64):

well, subtract .5 from it as an adjustment

OpenStudy (anonymous):

0.2123?

OpenStudy (amistre64):

thats what im getting yes. i cant say thats the correct answer tho simply becuase i dont have the flow charts that tell us when to sqrt(n) and when not to ...

OpenStudy (amistre64):

http://www.statisticshowto.com/when-to-use-a-t-score-vs-z-score/ this gives a little information

jimthompson5910 (jim_thompson5910):

We're dealing with the xbar distribution (aka the distribution of sample means) because it says "What is the probability that the `average` of a random sample of examination scores of 25 students will be between 450 and 500?" keyword: average this distribution has mean of mu and standard deviation of sigma/sqrt(n)

OpenStudy (amistre64):

so we might have to divide .56 by 5 ....

OpenStudy (anonymous):

Column 0.6 and row 0.1?

jimthompson5910 (jim_thompson5910):

\[\LARGE z = \frac{x-\mu}{\frac{\sigma}{\sqrt{n}}}\] \[\LARGE z = \frac{x-\mu}{\sigma}*\sqrt{n}\] If you already calculated \(\LARGE \frac{x-\mu}{\sigma}\), then you just need to multiply that result by sqrt(n) = sqrt(25)=5